OpenStudy (anonymous):

State whether the given measurements determine zero, one, or two triangles. A = 80°, a = 24, b = 50

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

zero

OpenStudy (anonymous):

how'd u get thaat?!

OpenStudy (loser66):

\[\frac{sinA}{a}=\frac{sinB}{b}\\sinB= \frac{b sinA}{a}\\sin B = 0.89 \]which leads to B = 197, That is not angle in a triangle which total degree is 180

OpenStudy (anonymous):

ahhh you're right :D thank you!!

OpenStudy (anonymous):

do you mind helping me with another?>

OpenStudy (loser66):

I have something to do, cannot stay. I am sorry, just post. There is someone helps you.

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

wrong